### v1

# In-vitro measurements coupled with in-silico simulations for stochastic calibration and uncertainty quantification of the mechanical response of biological materials

Abstract
In the standard Bayesian framework, a likelihood function is required, which can be difficult or computationally expensive to derive for non-homogeneous biological materials under complex mechanical loading conditions. Here, we propose a simple and practical approach based on in-vitro measurements and in-silico simulations using the likelihood-free Bayesian inference with the finite element method simultaneously for stochastic calibration and uncertainty quantification of the mechanical response of biological materials. We implement this approach for distal, middle, and proximal human Achilles tendon specimens obtained from diabetic patients post-amputation. A wide range of in-vitro loading conditions are considered, including one-step and two-step relaxation, as well as incremental cyclic loading tests. In-silico simulations are performed for the tendons assuming a fiber-reinforced viscoelastic response, which is modeled separately for the ground matrix and fiber components. Initially, the calibration of the specimen-specific parameters is predicted using Bayesian optimization and the sensitivity of each parameter is evaluated using the Sobol index and random forest. Then, the bounds of these parameters are used as priors, uniformly sampled, and coupled with in-vitro data in simulation-based approximate Bayesian computations to calibrate and quantify the uncertainty parameters for three loading cases. The results demonstrate that increasing the number of loading scenarios improves the accuracy of the specimen-specific best-fit parameter set. Furthermore, approximate Bayesian inference shows that in-silico simulations using the posterior parameters, can capture the uncertainty bounds of in-vitro measurements. This approach provides a useful framework for stochastic calibration of constitutive material model parameters without the need to derive a likelihood function, regardless of the specimen’s geometry or loading conditions.
Keywords: In-vitro measurements, in-silico simulations, biological materials, uncertainty quantification, stochastic material calibration, approximate Bayesian computation, finite element method
1 Introduction
In-silico mechanical modeling and simulation of biological materials has the potential to facilitate advancements in every aspect of the medical field. It offers numerous advantages, such as a promising tool for pre-clinical testing of medical devices, aiding in the diagnosis and evaluation of targeted treatments for patient-specific applications, and supplementing experimental investigations when in-vivo experiments are not possible [viceconti_possible_2021, smit_silico_2024, VV40, pekedis_2025_2]. It can also model patient-specific tissue or organ-level behavior and, more generally, account for variations within a population. One such factor that significantly influences the in-silico simulation outcome is the constitutive material model, as well as the parameters used in it. The model and its calibrated biomechanical parameters should accurately reflect and mimic the tissue’s behavior. These parameters are typically calibrated using a non-linear least squares fit of the model to mechanical testing data from experiments such as tension [ogden_fitting_2004, pekedis_2021], compression [budday_fifty_2020], and biaxial testing [niestrawska_microstructure_2016]. For most calibration procedures, the choice of a constitutive model is assessed by the quality of fit (e.g., $R^{2}$ and the value of the root mean square error RMSE). This technique is efficient and powerful for patient or specimen-specific calibration. However, because most biological materials are nonhomogeneous and anisotropic, deploying experimental tests using standard mechanical tests may not be enough for mechanical characterization of the tissue. Therefore, calibration of the model parameters directly from the experimental data may not always be possible or relevant. In some studies, the inverse technique is performed to overcome these issues [MENGONI2021105, narayanan_inverse_2021]. In brief, the inverse approach is considered as the estimation of input parameters (material properties, geometric features, loading and boundary conditions, etc.) from output observations (displacement, strains, velocity). Once a forward model is created, the input parameters are predicted using experimental observations with an optimization technique. This method is powerful for patient/specimen-specific material calibration when inter-variability is not considered.
In addition to inhomogeneous characteristics, biological materials often exhibit intrinsic variability in their biomechanical response when subjected to external loads. The inter-subject or inter-specimen variability and uncertainty in biomechanical properties of these materials are high, which may be influenced by many factors such as the donor’s age, genetics, the tissue’s pathological state, gender, and other specific factors related to the tissue characteristics. For in-vitro samples, embalming medium, temperature, humidity and duration are the additional factors that can affect the biomechanical properties. Oftentimes, in-silico model is hindered not only by the incompleteness of the theoretical approach, but also by uncertainties due to mentioned factors that arise in the experimental observations. Thus, there is an urgent need to ensure that it can capture the intrinsic population of variability and not only representative of ‘mean’ individuals.
Uncertainty can also arise due to a lack of knowledge of the physical model of interest (insufficient initial and boundary conditions), as an addendum to the inherent variation in material properties [oberkamf, henninger_validation_2010]. However, most of the studies implemented in the field of biomechanics have mainly investigated uncertainties due to variabilities in material properties using the Bayesian framework [mihai_LA, laz]. In this framework, once the prior distribution of the parameters, in conjunction with a likelihood function, is constructed, the posterior distribution is determined using random sampling methods. There has been much research in the field of brain mechanics [madireddy_bayesian_2015, teferra_bayesian_2019, brewick_uncertainty_2018, staber_stochastic_2017, mihai_LA]. However, the standard Bayesian framework requires a likelihood function, which can be difficult or computationally expensive to derive for nonhomogeneous materials under complex biomechanical loading conditions. To overcome on these such issues, an approximate Bayesian computation (ABC) approach that does not use the likelihood function has been proposed [toni_simulation-based_2010, schalte_pyabc_2022] and implemented recently in some fields including structural health monitoring [Guofeng] medical decision making [Huaqing] and operation research management [garud]. However, the coupling of in-silico simulations for stochastic calibration and uncertainty quantification using likelihood-free Bayesian computation in the field of mechanics of biological materials has been largely unaddressed to date, to the best of our knowledge. The most attractive aspect of this approach is that the likelihood function is not required to be explicitly represented in a closed analytical form. Once an in-silico model is built to mimic the in-vitro conditions, it directly bypasses the need for explicit likelihood evaluation and relies on comparing the simulated data from the in-silico model with experimental observations from the in-vitro test. This approach can be implemented in a wide range of in-vivo and in-vitro applications, ranging from predicting the mechanical characteristics change of tissue due to the risk of pathogenic disease to planning surgical procedures. We note that this study is not targeted to develop or present a constitutive material model for connective tissue, but rather to show how likelihood-free inference can be used directly with the numerical simulator and measurements. Any suggested model can be directly embedded in the simulator.
In this work, we suggest a simple approach based on in-vitro measurements, in-silico simulations, and approximate Bayesian computation for stochastic calibration and uncertainty quantification of material parameters. The paper is structured as follows. First, a specimen-specific calibration procedure using Bayesian optimization is presented. Next, the sensitivity analysis and stochastic calibration steps implemented in ABC are detailed in the following section. Finally, the results obtained from the methodology are presented and summarized, together with some concluding remarks.
2 Materials and Methods
2.1 In-vitro measurements
The approach is performed on recently published data of Achilles tendons from diabetic patients who underwent below or above knee amputation [pekedis_2025]. In summary, the tendons were harvested from the proximal end to the insertion of the Achilles tendon at the calcaneus. Their length was about 120 to 130 cm and were partitioned into three equal sections (distal, middle, and proximal site). Then, they were resectioned to form a standard dog bone specimen. The dimensions of each specimen at the measurement site had a gauge length of approximately 10 mm, a width of $\sim$ 8 mm and a thickness of $\sim$ 4 mm. Uniaxial displacement-type loading was performed to collect force measurements. First, samples were preloaded at 0.1 MPa to achieve parallel alignment of the fibrils, and then preconditioned for 10 cycles with a strain amplitude of 2% strain at a strain rate of 0.5% $s^{-1}$. The test protocols were as follows:
  • One-step relaxation: Initiates with a ramp strain of 0.6% ${s^{-1}}$ followed by a constant strain of 4% for a relaxation duration of 200 s (Figure 1a).
  • Two-step relaxation: Two consecutive steps of 2% strain (ramp strain of 0.6% $s^{-1}$) followed by a 200-s relaxation time (Figure 1b).
  • Incremental cyclic: Incremental cyclic loading ranging from a strain amplitude of 2% to 4.25% with an increment of 0. 2% strain in each cycle (Figure 1c).
Further details of the tests and in-vitro measurements are reported in previous work [pekedis_2025]. A typical stress-strain observation of a specimen for three loading scenarios is presented in Figure 1. The force and displacement signals have a sampling rate of 100 Hz. In-vitro measurements include 81 datasets (9 specimens x 3 sites x 3 test protocols). Each of these data sets consists of three dimensions: the first is time, the second is displacement, and the third is force. The data sets are organized and saved as PYTHON pickle for further serialization processes between in-vitro measurements and in-silico simulations for sensitivity, optimization and Bayesian inference tasks.
Figure 1: Typical experimental in-vitro measurements observed for a specimen. (a) One-step relaxation. (b) Two-step relaxation. (c) Incremental cyclic loading
2.2 Fiber reinforced visco elastic model
Soft biological tissues, such as the Achilles tendon, are collagenous tissues that connect the calf muscle to the calcaneus. Their primary function is to transmit tensile forces and act as mechanical dampers. These tissues are primarily composed of water, Type I collagen fibers, and proteoglycans. They can be viewed as fibrous composites, consisting of a ground matrix embedded with families of collagen fibers. The tendon is modeled using the fiber-reinforced viscoelastic (FRVE) approach, in which the solid matrix is divided into a nonfibrillar part representing the proteoglycan matrix and a fibrillar part representing the collagen fibers. The biomechanical response of these constituents, which drives the approximate Bayesian computation (ABC), is simulated using the FEBio framework [maas_febio_2012]. In this FRVE model, two solid mixtures are considered: the ground matrix, representing the proteoglycan matrix, and the collagen fibrils, representing the fibers. Viscoelastic effects are incorporated for both constituents. Specifically, the ground matrix is modeled as a compressible visco linear elastic material, whereas the collagen fibrils are assumed to be compressible visco nonlinear elastic materials that respond only to tensile loads. The relaxation function for the fibril and matrix components is given by:
$G(t)=g_{0}+\sum_{i=1}^{n}{g_{i}e^{-\frac{t}{\tau_{i}}}}$ (1)
where $t$ is time, $g_{i}$ and $\tau_{i}$ are $i$’th viscoelastic characteristic coefficients. Additionally, the strain energy density for the fiber component is follow:
$\psi_{n}=\begin{cases}0<I_{n},&0,\ \text{under compression}\\ \frac{\xi}{\alpha\beta}(\exp[\alpha(I_{n}-1)^{\beta}]-1),&1\leq I_{n}\leq I_{0},\text{for nonlinear toe region}\\ B(I_{n}-I_{0})-E_{f}(I_{n}^{1/2}-I_{0}^{1/2})+\frac{\xi}{\alpha\beta}(\exp[\alpha(I_{0}-1)^{\beta}]-1),&I_{n}\geq I_{o},\ \text{for linear region }\end{cases}$ (2)
where $E_{f}$ is the fiber modulus in the linear region, $\alpha$ is the coefficient of exponential component, $\beta$ is the power-law component in the toe region, $I_{0}=\lambda_{0}^{2}$ and $I_{n}=\lambda_{n}^{2}$. Here, $\lambda_{n}$ represents the current stretch ratio, and $\lambda_{0}$ denotes the stretch ratio in the transition from the toe region to the linear region. Additionally, $\xi$ and $B$ are as follows:
$\xi=\frac{E_{f}(I_{0}-1)^{2-\beta}\exp[-\alpha(I_{0}-1)^{\beta}]}{4I_{0}^{3/2}[\beta-1+\alpha\beta(I_{0}-1)^{\beta}]}$ (3)
$B=E_{f}\frac{2I_{0}[\beta-0.5+\alpha\beta(I_{0}-1)^{\beta}]-1}{4I_{0}^{3/2}[\beta-1+\alpha\beta(I_{0}-1)^{\beta}]}$ (4)
In-vitro measurements showed that $\lambda_{0}$ is approximately 1.01 and is assumed to be a fixed parameter. The above equation indicates that the fiber modulus reduces to zero unless $\beta=2$. Thus, we set $\beta=2$ in the in-silico simulation. In summary, the components constituting the matrix are assumed to exhibit viscoelastic behavior, and the fiber is modeled as a fiber-reinforced viscoelastic material. The descriptions of parameters, the search ranges for Bayesian optimization, and their priors are summarized in Table  1.
Table 1: Parameters used in Bayesian optimization (BO) and approximate Bayesian computation (ABC) for the three sites.
Constituent Parameter Description Ranges (Min, Max)
Fiber Ef (MPa) Fiber modulus (160, 275)
a (-) Exponential argument (0, 50E3)* and (0, 10E3)**
g1f (MPa) 1st viscoelastic characteristic (0.8, 1)
g2f (MPa) 2nd viscoelastic characteristic (0.8, 2.1)
g3f (MPa) 3rd viscoelastic characteristic (0.8, 2.1)
$\tau_{1\text{f}}$ (s) 1st viscoelastic relaxation (0.1, 2)
$\tau_{2\text{f}}$ (s) 2nd viscoelastic relaxation (2, 20)
$\tau_{3\text{f}}$ (s) 3rd viscoelastic relaxation (20, 200)
Ground Matrix Em (MPa) Elastic modulus (11, 70)
g1m (MPa) 1st viscoelastic characteristic (0.8, 1.5)
g2m (MPa) 2nd viscoelastic characteristic (0.8, 2.1)
g3m (MPa) 3rd viscoelastic characteristic (0.8, 2.1)
$\tau_{1\text{m}}$ (s) 1st viscoelastic relaxation (0.1, 2)
$\tau_{2\text{m}}$ (s) 2nd viscoelastic relaxation (3, 20)
$\tau_{3\text{m}}$ (s) 3rd viscoelastic relaxation (20, 200)
$v_{\text{m}}$ (-) Poisson’s ratio (0, 0.4)
Min and Max values represent the search ranges used in BO.
Min and Max values are also used as uniformly distributed prior ranges for ABC.
*represents $\alpha$ search ranges in BO and, **denotes uniformly distributed prior ranges of $\alpha$ for ABC.
2.3 In-silico simulations
Numerical methods such as the finite element method, meshless methods, or the finite volume method can be used as simulators to bridge the gap between in-vitro measurements and Bayesian computation for stochastic calibration and quantification of uncertainty in the biomechanical response of biological materials. Here, we used the finite element method as a simulator. The geometric dimensions of the in-vitro sample are approximately $10\times 8\times 4\,\text{mm}$ (gauge length $\times$ width $\times$ thickness).
To investigate the effectiveness of the approach and reduce the CPU solution time, we modeled half of the sample by applying a symmetric boundary condition using a single hexagonal element. As stated in the previous section, we considered an FRVE material property for the tendon (Figure 2). The boundary conditions of the in-silico model are similar to the loading conditions of the in-vitro test: the nodes located in the lower section are restricted only in the vertical direction ($u_{z}(t)=0$), and the tensile displacement observed from the in-vitro measurement ($u_{z}(t)$) for the three loading tests is applied longitudinally to the upper four nodes. A symmetric boundary condition is applied to the four nodes on the left side ($u_{y}(t)=0$) (Figure 2). We note that additional in-silico outputs, such as the deformation gradient, the strain distribution, or the direct stress, can also be extracted and used with ABC if the corresponding measurements are available. Typically, full-field deformation measurements can be collected via image-based techniques, such as digital image correlation. We believe that higher-dimensional data from in-vitro observations will lead to more sensitive calibrated parameters.
Figure 2: Specimen-specific parameter calibration steps using in-vitro measurements, in-silico simulations, and Bayesian optimization
2.4 Bayesian optimization and sensitivity analysis
To determine specimen-specific biomechanical parameters, and use the minimum and maximum bounds of these parameters for uncertainty quantification scheme, Bayesian optimization (BO) is implemented using scikit-optimize library. BO is an iterative framework focused on optimizing an expensive and unknown objective function $f(x)$, where $x\in\mathbb{R}^{d}$ is a vector of $d$ parameters. The goal is to solve
$\mathbf{x}^{*}=\arg\min_{\mathbf{x}\in\mathcal{A}}f(x)$ (5)
where $\mathcal{A}$ is the search space. The gradient information is unavailable for $f(x)$ and is considered a “black box” function. The core of BO is based on constructing a probabilistic surrogate model, a Gaussian process (GP) to approximate $f(x)$. It models the unknown objective function $f(x)$ by providing a posterior distribution given the observed data. With its simplest form, the resulting prior of GP is written as:
$f(x)\sim\mathcal{GP}(\mu(x),k(x,x^{\prime}))$ (6)
where $\mu(x)$ is the mean function, showing the prior belief about the objective function and $k(x,x^{\prime})$ is the covariance (or kernel) function. Once building the surrogate model, Bayesian optimization selects the next point $x_{\text{next}}$ to evaluate by maximizing an acquisition function. A common acquisition function used in scikit-optimize is Expected Improvement (EI) and defined as:
$\text{EI}(x)=\mathbb{E}\left[\max(f^{*}-f(x),0)\right]$ (7)
where $f^{*}$ is the current best observed value of $f(x)$. The $f(x)$ function at any $x$ point is normally distributed with mean $\mu(x)$ and variance $\sigma^{2}(x)$, and the improvement is modeled as:
$\text{EI}(x)=(\mu(x)-f^{*})\Phi\left(\frac{\mu(x)-f^{*}}{\sigma(x)}\right)+\sigma(x)\phi\left(\frac{\mu(x)-f^{*}}{\sigma(x)}\right)$ (8)
where $\Phi(\cdot)$ is the cumulative distribution function (CDF) and $\phi(\cdot)$ is the probability density function (PDF) of the standard normal distribution, respectively. This acquisition function seeks regions where the model predicts the largest improvement over the current best value. To find the next sampling point $x_{\text{next}}$, the acquisition function $\text{EI}(x)$ is maximized. Once a new sample $x_{\text{next}}$ is selected and the objective function $f(x_{\text{next}})$ is evaluated, the surrogate model (the Gaussian Process) is updated. The process continues until a stopping criterion is met. Further details of BO are documented in literature [frazier2018tutorialbayesianoptimization]. We have made a couple of trials, and observed that a stop criteria with 100 fixed number iterations is reasonable for tuning the biomechanical parameters. We note that a higher number of iterations increased the computational cost. The objective function to be minimized is defined as the summation of the root mean square error (RMSE) between experimental (in-vitro) and numerical (in-silico) total reaction force values for the three loading scenarios. It is given by:
[…]
where $R_{ij}$ and $\mathcal{S}{(u_{ij},x})$ are the experimental and numerical reaction forces, respectively, for the parameter set $x$ at displacement $u_{ij}$, and data point i. The term $n_{j}$ denotes the number of data points for test loading j, where j = 1 corresponds to the one-step relaxation, j = 2 to the two-step relaxation, and j = 3 to the incremental cyclic loading scenario. The factor of 2 in the equation arises from the symmetric boundary conditions modeled in the finite element (FE) simulation. Simply, the $(R_{ij},u_{ij})$ are in-vitro observation pairs, and $\mathcal{S}(u_{ij},x)$ is the unknown force function that extracts the corresponding data from the FE simulator. The BO tries to find the model parameters $x$ that minimize the function $f(x)$ by repeatedly evaluating the function $\text{EI}(x)$, which calls the simulator (FEBio) to solve the black-box problem. The parameter search ranges used during BO are given in Table  1, and the initial search ranges for these parameters are predicted as follows. First, the viscoelastic characteristics $g_{i}$ and $\tau_{i}$ are estimated by fitting on some trial one-step relaxation measurements using the Levenberg-Marquardt method. Then, the bounds of the estimated parameters are extended and used as search ranges for BO. Similar search ranges are considered for the fiber and matrix components. For the approximate prediction of the modulus, we initially evaluated it by analyzing the linear region of a typical incremental cyclic test loading measurement and found that the modulus of elasticity for the composite tendon (matrix + fiber) is around 350 MP. Although, to the best of our knowledge, no study has experimentally quantified the elastic modulus of the fiber and elastic component in the human Achilles tendon separately, it has been stated in some rat studies that the fiber-to-matrix ratio is between 5 and 10. [dutov_measurement_2016, khayyeri_fibre-reinforced_2015, lewis_modeling_1997]. We considered these uncertain ranges during the determination of the search ranges.
Sensitivity and feature importance analyses are performed to investigate the influence of each parameter on the FRVE model for each test loading scenario. Feature importance is determined using the random forest algorithm implemented in scikit-learn, which is based on impurity reduction. The importance of a feature is computed as the total reduction of the criterion observed by the feature, known as the Gini importance [breiman_classificatin]. The importance of a feature $I(x_{i})$ is given by:
$I(x_{i})=\frac{1}{n_{t}}\sum_{1}^{n_{t}}\sum_{s\in S_{t}(x_{i})}\Delta I_{s}$ (10)
where $n_{t}$ is the number of trees in the forest, $S_{t}(x_{i})$ is the set of all splits for the feature $x_{i}$ in the tree $t$, and $\Delta I_{s}$ is the decrease in the splitting criteria at split $s$. The final importance values are then normalized to 100 over all features. More details on the random forest algorithm are documented in the literature [breiman_classificatin].
As an addendum to the feature importance analysis, we implemented Sobol sensitivity analysis to investigate the sensitivity of the input parameters on the model output using salib [herman_salib_2017]. It is a variance-based approach to quantify the parameter’s sensitivity as well as their contribution to the model. Basically, it decomposes the total variance of the output $Y$ into contributions that are attributable to input parameters and their interactions. Let us assume that $Y=f(X_{1},X_{2},\dots,X_{d})$ represents the model output as a function of $d$ input parameters $X_{1},X_{2},\dots,X_{d}$. The total variance of $Y$ is given by:
[…]
where $p(\mathbf{X})$ shows the joint probability density function for the inputs. The variance can be decomposed as:
$V=\sum_{i=1}^{d}V_{i}+\sum_{1\leq i<j\leq d}V_{ij}+\cdots+V_{12\ldots d}$ (12)
where $V_{i}=\mathrm{Var}\left(E[Y\mid X_{i}]\right)$ represents the main effect of input $X_{i}$, and $V_{ij}=\mathrm{Var}\left(E[Y\mid X_{i},X_{j}]\right)-V_{i}-V_{j}$ shows the interaction effect between $X_{i}$ and $X_{j}$. Higher-order terms denote interactions between more input parameters. The Sobol indices are dimensionless quantities that measure the relative contribution of each term to the total variance. The first order sensitivity index is given by:
$S_{i}=\frac{V_{i}}{V}$ (13)
The 1st order index shows the proportion of variance that is attributed to $X_{i}$. Total-order sensitivity index is given by:
$S_{T_{i}}=1-\frac{V_{\sim i}}{V}$ (14)
where $V_{\sim i}$ is the variance of the output excluding $X_{i}$. The total-order index accounts for both the main effect and all interaction effects involving $X_{i}$. The interaction terms (e.g., $S_{ij}$) can be calculated similarly by dividing the corresponding variance term by the total variance. However, the sensitivity due to interaction is not covered in this study. We implemented Saltelli sampling to create the parameter space of 20,000 points for each of the 16 input parameters of the FRVE model. The outputs are the results computed from in-silico simulations. The sampling ranges used for BO and ABC are also used during Saltelli sampling (Table  1). We used the minimization function values given in Eq. 9 for j = 1, 2, 3, as the results of the Sobol sensitivity analysis. More details on the Sobol sensitivity analysis are reported in the previous study [SALTELLI2010259].
2.5 Approximate Bayesian Computation
Although the standard Bayesian framework allows for updating prior beliefs about parameters based on the likelihood of the data given those parameters, evaluating the likelihood $P(\mathbf{y}_{\text{obs}}|\boldsymbol{\theta})$ is often infeasible, specifically for complex models, but the model can generate synthetic data $\mathbf{y}_{\text{sim}}$ given parameters $\boldsymbol{{\theta}}$. We implemented likelihood-free approximate Bayesian computation (ABC) for stochastic calibration and uncertainty quantification of posterior parameters using PyABC [schalte_pyabc_2022]. The framework is based on an executable “black-box” forward process model that simulates data using given model parameters [Price02012018, wilkonson_2013]. In ABC, the computation of the likelihood is replaced by the comparison between observed (in-vitro measurements) $\mathbf{y}_{\text{obs}}$ and simulated data (in-silico simulations) $\mathbf{y}_{\text{sim}}$. Given the prior distribution of $P(\boldsymbol{\theta})$ of parameters $\boldsymbol{\theta}$, the target is to approximate the posterior distribution $P(\boldsymbol{\theta}|\mathbf{y}_{\text{obs}})$ given observed data $\mathbf{y}_{\text{obs}}$, a model $f(\mathbf{y}_{\text{sim}}|\boldsymbol{\theta})$, and a prior distribution $P(\boldsymbol{\theta})$. Typically, it is based on generating simulated data $\mathbf{y}_{\text{sim}}$ for parameters $\boldsymbol{\theta}$ drawn from the prior distribution and retaining parameters for which $\mathbf{y}_{\text{sim}}$ is sufficiently close to $\mathbf{y}_{\text{obs}}$. The posterior is approximated as:
$P(\boldsymbol{\theta}|\mathbf{y}_{\text{obs}})\approx P(\boldsymbol{\theta})\cdot\mathbb{I}(d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\leq\epsilon)$ (15)
where $d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})$ is a distance metric, $\epsilon$ is a tolerance threshold, and $\mathbb{I}(\cdot)$ is an indicator function that equals 1 when the condition is satisfied $d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\leq\epsilon$ and 0 otherwise $d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\geq\epsilon$. The output of the algorithm is a sample set of parameters from the distribution $P(\theta|d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})\leq\epsilon)$. Direct ABC implementation can be inefficient, particularly for small ϵ, as it reduces the number of accepted parameters. To overcome this, pyABC performs a Sequential Monte Carlo (SMC) technique that progressively refines the posterior estimation. It facilitates robust and efficient inference for a broad spectrum of applications. Further detail of ABC model is documented in previous studies [toni_simulation-based_2010].
In this study, the prior distributions for the parameters are defined based on a uniform distribution (Table  1). The distance between observed and simulated data is computed using the Euclidean distance (L2) metric. To refine the parameter estimates, the $\epsilon$ tolerance was adaptively adjusted in the ABC SMC process. The initial tolerance threshold of $\epsilon$ is set as 0.3. We note that we selected these options by running some trials and focusing on where the simulated data $\mathbf{y}_{\text{sim}}$ most closely matched the observed data $\mathbf{y}_{\text{obs}}$
2.6 Implementation procedures
The implementation begins with specimen-specific parameter calibration to determine the approximate values of the FRVE parameter set using an optimization technique. We employed Bayesian Optimization (BO), and the procedures and stages of its implementation are illustrated in Figure  2, where the numbers indicate data processing steps. Initially, in-vitro test samples are harvested and re-sectioned (1). Then in vitro test loading protocols are applied and their measurements are acquired, categorized and saved as a serialized dataset file (2). As a preliminary study, we chose a simple uniaxial tensile test for the three test loading scenarios. For each test, the reaction force $R(t)$ is obtained as a function of displacement $u(t)$. This process is applied to all samples. Next, the search space is defined, and the objective function is formulated (3). In-silico simulations using the finite element method are performed to obtain time vs. reaction force values (4). The objective function is computed as the difference between in-silico and in-vitro reaction force observations (5). Once the objective function is minimized in the current step, the next target is to optimize the acquisition function, balancing exploration and exploitation (6-7). The model is then updated by incorporating a new observation and retraining the surrogate model. Finally, steps 3-7 are iterated until the stopping criteria are met (8). To reduce computational time, we set the stopping criterion at 100 iterations, as this was sufficient for calibrating specimen-specific parameters. The implementation of in-vitro measurements, coupled with in-silico simulation and ABC, resembles the BO stages at the pre-processing steps. However, the main difference is that instead of evaluating specimen-specific observations, ABC includes all specimens’ measurements, allowing for the stochastic calibration of parameters, including uncertainty. Further details are as follows (Figure 3).
  1. 1.
    In-vitro test samples are harvested and re-sectioned into an appropriate shape for the biomechanical test. A higher number of samples allows for better quantification.
  2. 2.
    In-vitro biomechanical tests are performed to obtain output observations. Here, the resulting reaction forces $R(t)$ are referred to as $\mathbf{y}_{\text{obs}}$. As an addendum to the one-step and two-step relaxation loadings, we include only single-cycle observations (loading-unloading) from the incremental cyclic test during ABC inference.
  3. 3.
    The importance and sensitivity of each parameter are determined using the Random Forest and Sobol methods. Once the minimum and maximum bounds obtained from BO across all samples are determined, these wide bounds are used as priors in ABC. Initially, parameters $\boldsymbol{\theta}$ are drawn from the prior uniform distribution as: $\boldsymbol{\theta}\sim P(\boldsymbol{\theta})$.
  4. 4.
    The next step is to perform forward simulations to drive ABC. We implemented FEBio as the simulator and created a template that includes the FRVE material model and boundary conditions that mimic the experiment. FEBio input files (.feb) are xml type files, and we used the jinja package to create the FEBio input template file and serialize the data between ABC and the simulator. Once ABC determines the input parameter set $\boldsymbol{\theta}$, it is passed directly to this simulator template, and the FEBio simulator generates simulated data $f(\mathbf{y}|\boldsymbol{\theta})$ using $\boldsymbol{\theta}$ as: $\mathbf{y}_{\text{sim}}\sim f(\mathbf{y}|\boldsymbol{\theta})$.
  5. 5.
    The distance between the observed and simulated data is computed as: $d=d(\mathbf{y}_{\text{obs}},\mathbf{y}_{\text{sim}})$, where $d(\cdot)$ is a Euclidean distance metric.
  6. 6.
    The parameter $\boldsymbol{\theta}$ is retained if the distance satisfies: $d\leq\epsilon$; otherwise, return to step 4.
  7. 7.
    The joint plots of the posterior parameters are visualized, and the lower, mean and upper bounds, as well as the confidence interval 95% of the inferred simulation $\mathbf{y}_{\text{sim}}$, are plotted.
Figure 3: Stochastic parameter calibration steps using in-vitro measurements, in-silico simulations and approximate Bayesian computation
3 Results
3.1 Bayesian optimization
We first performed specimen-specific parameter calibration using Bayesian optimization (BO), and the results observed for a typical specimen are represented in Figures 4, 5, 6, and 7. Figure 4 shows that the minimum $f(x)$ function converges rapidly, even with only a few initial iterations. This characteristic makes BO popular in machine learning for hyperparameter optimization, where the learning algorithm is treated as a black-box function. Consider the convergence plot given in Figure 4, although the initial minimum value $f(x)$ is 65 N, after 100 iterations (or function calls), it decreases to less than 20 N. In addition, Figure 5 presents the in-vitro measurements obtained from experiments and the in-silico observations generated using optimized FRVE parameter sets in BO. Here, we note that the terms ‘ground truth’, ‘in-vitro measurements’, and ‘experimental observations’ are often used interchangeably to convey a similar meaning. Furthermore, ‘in-silico simulation’, ‘black-box simulator’, and ‘numerical observations’ are also often used interchangeably to refer to a similar meaning. Figure 5 shows that the minimum $f(x)$ values of the objective function, determined by incremental cyclic loading test measurements in one step, two steps, and incremental steps, are 5.693 N, 3.871 N, and 13.232 N, respectively. These results show that the FRVE model can capture the mechanical response of the tendon. Furthermore, we have also included the updated parameter set that was sampled during BO for each iteration as additional figures in this study (See supplementary Graphs G1).
We performed the BO process for 100 calls of the black-box simulator, and the parameter sensitivity can probably be improved once the iteration number is increased. However, this can increase computational time, and we have not focused on characterizing the parameters with a high rank of sensitivity. The target of using BO is to determine how the specimen-specific parameters vary across specimens and to use these parameters to predict the priors for ABC inference. We present some parameters for matrix and fiber constituent pairs that were sampled during BO updating as scatter (Figure 6) and contour (Figure 7) plots to visualize where they are mainly clustered. For example, consider the sample points of $E_{m}$ and $E_{f}$; the elastic modulus of the matrix and fiber components is predominantly updated around 60 MPa and 240 MPa, respectively. The remaining parameters for a typical specimen-specific calibration are given as additional figures (See Supplementary Figure S1 and S2). We then automated the process to characterize the remaining 26 samples, and the overall results are given in Table 2. These results suggest that inter-subject or inter-specimen variability and uncertainty in the biomechanical properties of these materials are high.
Figure 4: Typical convergence plot of a specimen during Bayesian optimization
Figure 5: In-vitro (ground truth) and in-silico (simulation) observations of a typical sample with parameters tuned with Bayesian optimization.
Figure 6: Visualization of the sample points for some parameters of a typical specimen during Bayesian optimization. The red star indicates the best-found parameters.)
Figure 7: The influence of each search space on the objective function for some parameters observed in a typical specimen during Bayesian optimization. The red star indicates the best-found parameters
Table 2: FRVE model parameters tuned with Bayesian optimization
Site ID $\tau_{\text{1m}}$ (s) $\tau_{{2\text{m}}}$ (s) $\tau_{{3\text{m}}}$ (s) g1m (MPa) g2m (MPa) g3m (MPa) Em (MPa) vm $\tau_{\text{1f}}$ (s) $\tau_{\text{2f}}$ (s) $\tau_{\text{3f}}$ (s) g1f (MPa) g2f (MPa) g3f (MPa) Ef (MPa) $\alpha$f RMSE (N)
Distal 1 0.578 18.626 197.901 1.468 2.100 0.800 59.900 0.240 0.386 17.898 51.569 0.957 1.575 0.917 225.145 2.05E3 22.887
2 0.407 4.937 60.607 1.315 1.760 1.944 70.000 0.254 1.644 15.729 83.984 0.943 1.629 2.072 206.852 1.09E3 21.352
3 0.156 14.878 152.518 1.463 0.812 0.981 62.716 0.300 1.549 8.717 20.000 0.905 1.155 0.800 216.049 1.30E3 20.892
4 1.138 17.810 146.273 1.461 1.715 2.100 31.571 0.240 2.000 20.000 133.023 1.000 2.100 2.100 203.115 0.00E0 26.799
5 0.819 9.548 194.489 1.352 1.933 0.800 58.204 0.121 2.000 10.558 40.376 0.905 1.245 2.100 244.672 8.62E2 24.957
6 0.905 18.524 23.824 1.500 1.795 0.800 34.364 0.001 2.000 18.802 169.058 0.821 2.100 2.100 167.979 2.23E2 21.311
7 2.000 11.041 142.566 1.500 2.100 2.100 70.000 0.400 2.000 20.000 200.000 1.000 2.100 2.100 229.886 0.00E0 14.737
8 2.000 3.000 113.385 1.472 1.602 2.100 70.000 0.228 2.000 20.000 20.000 0.913 2.100 2.100 182.576 0.00E0 25.667
9 2.000 20.000 20.000 1.500 2.100 2.100 46.866 0.000 2.000 20.000 200.000 1.000 2.100 2.100 167.365 0.00E0 17.196
Middle 1 2.000 3.000 20.000 0.800 1.678 2.052 63.561 0.400 1.347 20.000 196.344 0.952 1.973 2.100 160.000 0.00E0 21.989
2 0.921 3.616 153.302 0.800 1.162 2.038 70.000 0.00 0.954 20.000 200.000 0.890 2.100 2.100 195.264 2.86E4 29.958
3 1.548 6.478 20.000 1.500 0.817 2.100 22.412 0.352 2.000 20.000 164.040 0.912 2.100 2.100 201.470 1.04E4 9.382
4 2.000 20.000 200.000 1.500 2.100 2.100 70.000 0.400 2.000 20.000 199.514 1.000 2.100 2.100 164.599 0.00E0 41.498
5 1.970 13.086 69.866 1.168 1.259 2.049 11.000 0.091 1.689 20.000 200.000 0.800 2.100 2.100 199.684 0.00E0 23.908
6 2.000 3.000 200.000 1.500 2.100 2.100 36.661 0.400 2.000 20.000 200.000 0.986 2.100 2.100 183.405 0.00E0 27.924
7 2.000 20.000 200.000 1.405 2.100 2.100 45.681 0.400 2.000 20.000 131.544 1.000 2.100 2.100 160.000 0.00E0 36.654
8 2.000 15.718 200.000 1.500 2.100 2.100 27.279 0.248 1.920 20.000 200.000 1.000 2.100 2.100 183.405 0.00E0 15.228
9 1.116 16.968 125.238 1.500 0.800 2.096 11.000 0.026 0.789 6.431 113.803 0.800 2.100 2.100 188.961 2.61E2 19.570
Proximal 1 0.651 6.536 45.061 1.473 0.815 2.100 25.626 0.100 1.811 20.000 79.971 0.952 1.819 2.100 271.839 1.47E3 17.278
2 1.693 12.671 200.000 0.985 1.974 2.100 70.000 0.001 2.000 20.000 200.000 1.000 1.728 2.100 174.056 0.00E0 27.972
3 2.000 20.000 200.000 1.084 2.100 2.100 70.000 0.094 2.000 20.000 193.850 0.966 2.100 2.100 188.282 0.00E0 41.157
4 2.000 20.000 200.000 1.500 2.100 2.100 70.000 0.400 2.000 20.000 200.000 0.997 2.100 2.100 236.762 0.00E0 35.313
5 1.554 20.000 108.220 1.497 2.100 1.675 18.517 0.323 2.000 20.000 145.985 0.953 2.100 2.100 230.546 0.00E0 29.566
6 1.001 4.105 188.331 1.500 2.100 1.570 70.000 0.305 1.920 20.000 55.307 0.800 1.507 2.100 227.843 4.33E2 20.384
7 2.000 20.000 200.000 1.500 2.100 2.100 70.000 0.001 2.000 20.000 200.000 1.000 2.100 2.100 275.000 0.00E0 43.663
8 1.722 5.033 147.013 1.298 1.288 2.043 49.906 0.237 1.693 14.746 85.025 0.974 1.692 1.738 191.018 1.91E3 18.428
9 1.620 16.161 97.951 0.934 0.849 1.941 11.823 0.001 1.071 18.328 161.805 0.877 2.100 2.100 183.610 0.00E0 16.286
For the clinical aspect, correlation analysis between micro-scale matrix and fiber parameters calibrated via BO can also be performed. Here, to account for dependency due to repeated observations for the same patient’s specimen harvested from distal, middle, and proximal sites, we used the rm_corr() repeated measures correlation function from the Pingouin statistical package. Repeated measures reveal that the most significant correlations are observed between $\tau_{3m}$ and $\tau_{1f}$ (r = 0.8), $g_{3m}$ and $g_{3f}$ (r = 0.71), $\tau_{2m}$ and $\tau_{1f}$ (r = 0.69), $g_{2m}$ and $\tau_{1f}$ (r = 0.65), $g_{2m}$ and $g_{3f}$ (r = 0.64), while the only negative significant correlation exists between $E_{m}$ and $\tau_{1f}$ (r = 0.48) (p $<$ 0.05) (Figure 8).
Figure 8: Correlation levels between matrix and fiber variables. *Repeated measures correlation coefficient (r) is significant at the p $<$ 0.05 level
3.2 Sensitivity analysis
To determine the influence of each parameter on the objective function of FRVE model parameters, we investigate the percentage importance of the parameters using random forest and their sensitivity using Sobol. Here, we use the error function observations $\min f(x)$ given in Eq. 9 as outputs during sensitivity analysis and importance identification. The results of the random forest for a typical specimen in three loading conditions are represented in Figure 9. As expected, the elastic modulus of the fiber $E_{f}$, the exponential coefficient $\alpha_{f}$, and the elastic modulus of the matrix $E_{m}$ have the highest feature importance. Interestingly, the viscoelastic characteristics $g_{im}$ and relaxation time coefficients $\tau_{im}$ of the ground matrix at the $i$th Prony order are more important than the fiber’s parameters for the combined three-loading $\min f(x)$ function.
Figure 9: Variables feature importance using Random forest
For Sobol’s sensitivity analysis, the parameter space is created using Saltelli sampling. We generate 20,000 input sampling sets, each consisting of 16 parameters (8 for the ground matrix and 8 for the fiber), and collect the output $\min f(x)$ observations. Similar to the results from the random forest, $E_{f}$, $\alpha_{f}$, $g_{1m}$, and $E_{m}$ are the most sensitive parameters in the model. The total and first-order indices observed for the eight most sensitive parameters are given in Figure 10. Considering total-order index scores, the sequence from highest to lowest is as follows: $E_{f}$ (0.51), $\alpha_{f}$ (0.48), $E_{m}$ (0.45), $g_{1m}$ (0.36), $\tau_{1m}$ (0.34), $v_{m}$ (0.25), and $\tau_{1f}$ (0.22). These results help to focus on the most critical parameters when defining priors for simulation-based ABC inference.
Figure 10: Sensitivity of the parameters
3.3 Approximate Bayesian computation
We presented specimen-specific parameter identification results in the previous section and observed that the variability between samples is high and may be influenced by many factors, such as the donor’s age, genetics, the tissue’s pathological state, gender, and other specific factors related to tissue characteristics. To determine the uncertainty in these parameters and stochastically calibrate them, we implement likelihood-free inference in a probabilistic framework. The ABC inference is performed with an initial tolerance threshold of $\epsilon=0.3$ and a posterior sample of 40. The inferred parameters for the distal, middle, and proximal sites are given in Table 3. Consider $E_{f}$ at the distal site: the mean is 216.2 MPa, and the 95% HDI bounds are 164.6 MPa and 267.8 MPa.
Table 3: Posterior parameters inferred for the three sites
Parameter Distal Middle Proximal
Ef (MPa) 216.2 ± 31.12 (164.6, 267.8) 249.1 ± 46.47 (167.2, 324.9 238.5 ± 50.04 (161.4, 322.3)
$\alpha_{\text{f}}$ (-) 515.7 ± 224.4 (105.7, 883.0) 566.1 ± 275.9 (76.26, 962.6) 548.9 ± 278.9 (76.26, 962.6)
g1f (MPa) 1.305 ± 0.254 (0.813, 1.772) 1.288 ± 0.280 (0.881, 1.762) 1.252 ± 0.272 (0.881, 1.762)
g2f (MPa) 1.857 ± 0.602 (0.855, 2.828) 1.726 ± 0.676 (0.803, 2.789) 1.814 ± 0.652 (0.803, 2.789)
g3f (MPa) 1.754 ± 0.638 (1.001, 2.896) 1.849 ± 0.629 (0.847, 2.820) 1.741 ± 0.633 (0.847, 2.820)
$\tau_{\text{1f}}$ (s) 0.988 ± 0.559 (0.116, 2.007) 1.086 ± 0.661 (0.167, 2.077) 1.001 ± 0.623 (0.167, 2.077)
$\tau_{\text{2f}}$ (s) 11.31 ± 5.395 (2.347, 19.41) 13.06 ± 5.395 (2.798, 21.40) 12.79 ± 5.550 (3.023, 21.75)
$\tau_{\text{3f}}$ (s) 106.0 ± 49.25 (22.91, 189.5) 120.9 ± 57.21 (30.07, 218.5) 119.9 ± 55.14 (24.44, 218.5)
Em (MPa) 42.16 ± 14.74 (18.26, 67.07) 46.19 ± 18.76 (14.82, 77.67) 42.76 ± 18.36 (14.82, 79.59)
g1m (MPa) 1.499 ± 0.348 (0.941, 2.158) 1.626 ± 0.440 (0.929, 2.296) 1.600 ± 0.428 (0.929, 2.296)
g2m (MPa) 1.948 ± 0.537 (0.861, 2.742) 1.841 ± 0.600 (1.033, 2.830) 1.752 ± 0.582 (1.012, 2.830)
g3m (MPa) 1.809 ± 0.620 (0.885, 2.834) 1.946 ± 0.528 (1.059, 2.856) 1.898 ± 0.553 (0.896, 2.824)
$\tau_{\text{1m}}$ (s) 0.814 ± 0.447 (0.188, 1.671) 0.977 ± 0.479 (0.264, 2.042) 0.960 ± 0.531 (0.179, 2.042)
$\tau_{\text{2m}}$ (s) 12.99 ± 5.199 (3.993, 21.96) 12.12 ± 6.041 (3.521, 22.88) 12.40 ± 6.176 (3.152, 22.15)
$\tau_{\text{3m}}$ (s) 83.58 ± 46.82 (20.62, 192.6) 112.1 ± 55.61 (24.08, 212.1) 104.8 ± 53.14 (24.08, 212.1)
vm (-) 0.208 ± 0.095 (0.069, 0.392) 0.202 ± 0.121 (0.015, 0.399) 0.210 ± 0.118 (0.015, 0.399)
The values are represented as mean ± standard deviation.
The values given in round brackets are the highest density interval (HDI) bounds at 95% level
We have visualized the posterior joint distributions to investigate the relationships, correlations, and uncertainty between the parameters after updating beliefs based on in vitro measurement observations. Some parameter pairs plotted for the matrix and fiber components, as well as their distributions for the distal site, are shown in Figure 11. Consider $g_{1}$: there is a significant negative correlation between the matrix and the fiber component (r = -0.37, p $<$ 0.05). However, the remaining parameters shown in Figure 11 for the matrix ($\tau_{1m}$, $g_{2m}$, $E_{m}$, $v_{m}$) and the fiber ($\tau_{1f}$, $g_{1f}$, $g_{2f}$, $\alpha_{f}$) are not significantly correlated. The joint posterior plots inferred for 16 parameters for three sites are given in Supplementary Figures S3-S5. Once the calibration of the parameters is complete and their uncertain ranges for the highest density interval (HDI) 95% are inferred, the next step is to investigate how well the simulated data from the posterior $f(y_{sim}|\theta)$ can capture the uncertainties of in vitro observations $y_{obs}$. We perform forward simulations based on posterior parameters ($\theta$) for two cases. In the first case, we performed three simulations: one based on the mean and the other two based on the low and high bounds of $\theta$ given in Table 3. We then collected observations from $y_{sim}$. The inferred observation band is shown in cyan, and its bounds are represented in blue in Figure 12. In the second case, we performed 40 simulations, each corresponding to a single posterior parameter set, and then computed the 95% confidence interval (CI) of the 40 $y_{sim}$ observations. The observation band for the 95% CI and its bounds are represented in red (Figure 12). The results obtained in both cases demonstrate that the stochastically calibrated model parameters can capture the uncertainty band of in vitro measurements ($y_{obs}$). However, to characterize and simulate the soft tissue response under extreme uncertainty, it is preferable to use the 95% HDI bound of $\theta$, as demonstrated in case 1.
Figure 11: Posterior joint plots between same parameters of the matrix and fiber component for the distal site
Figure 12: In-vitro measurements (ground truth) and inferred simulation observations based on posterior samples for the three sites
In some clinical biomechanical studies, it may be necessary to compare the data obtained from different groups. This is usually carried out by applying parametric or non-parametric statistical tests to the results. This approach is effective within a frequentist framework. Although this study does not directly investigate the differences between the three sites, we performed unsupervised machine learning to determine whether the probabilistic data of the groups are clustered and separated. Here, we primarily used principal component analysis (PCA) and receiver operating characteristic (ROC) curves to investigate whether data from different groups could be effectively distinguished. Prior to PCA, we standardized the posteriors $\theta$ as $\frac{\theta-\mu}{\sigma}$, where $\mu$ is the mean and $\sigma$ is the standard deviation. Then, we processed 40 parameters with a 16-dimensional posterior parameter set through PCA to reduce them to a single-dimensional component using the scikit-learn library. The reduced component observations corresponding to a specific site were grouped in pairs and represented via ROC curves. A perfect separation between two groups has an area under the curve (AUC) of 1, while an AUC value around 0.5 indicates minimal separation. The results given in Figure  13 suggest that the posterior parameters are not well separated between the distal vs. middle (AUC = 0.50), the middle vs. proximal (AUC = 0.51) and the distal vs. proximal (AUC = 0.51) sites. Furthermore, the kernel density distribution exhibited comparable findings, suggesting a lack of clear separation (Figure 13). Finally, using stochastically inferred parameters from in-silico simulations, it is also possible to investigate the stress uncertainties of the fiber and matrix constituents. A typical stress band observed for the distal site is represented in Figure 14, and for middle and proximal results are given in Supplementary Figure S6.
Figure 13: ROC curves and kernel density estimation for the 1st PCA component
Figure 14: Inferred matrix and fiber stresses for the distal site
4 Discussion and Conclusions
In-silico simulations using computational techniques have significant potential to model very specific cases, ranging from patient-specific tissue behavior to a large variation in population. One significant factor that influences the outcomes of the simulation is the selected constitutive material model, as well as the parameters used in this model. Homogeneous isotropic hyperelastic materials can capture the biomechanical response of many deformable solids. However, most biological materials exhibit inherent variation in their response and may not be accurately represented by a constitutive model with a single set of deterministic parameters. In this approach, the model is typically verified through forward model evaluations using a certain number of experiments. For the backward problem, an inverse approach is mostly used to identify the input parameters from the output. This method is powerful for the characterization of patient- or specimen-specific materials. To account for the uncertainties in the experimental observations, the standard Bayesian approach is widely implemented in many studies [madireddy_bayesian_2015, teferra_bayesian_2019, brewick_uncertainty_2018, staber_stochastic_2017, mihai_LA]. However, it may not always be possible to derive and express the likelihood function in explicit form, particularly for complex loading scenarios. We suggest a likelihood-free approximate Bayesian computation based on a black-box simulator. We demonstrate its implementation for Achilles tendon specimens from diabetic patients, which include deep uncertainties in their in-vitro observations.
There are a wide variety of factors, including pathology, microstructural characteristics, tissue age, and embalming duration, that can significantly influence the biomechanical response. The goal is to stochastically calibrate parameters that represent the uncertain bounds of experimental observations without using the likelihood function. We consider the FRVE model for the representation of in-vitro simulations with 16 parameters. We employ ABC-free likelihood coupled with in-silico simulation and in-vitro measurement for the estimation of posterior parameters. Additionally, Sobol sensitivity analysis and feature importance through a random forest are performed to investigate the influence of parameters on the model. The overall results we observed are as follows. First, in vitro measurements exhibit extreme uncertainties and variability between samples is high. Second, Bayesian optimization converges rapidly with just a few initial iterations and calibrates 16 specimen-specific parameters. Third, sensitivity analysis and feature importance showed that $E_{f}$, $\alpha_{f}$, $g_{1m}$, and $E_{m}$ parameters have the highest influence on the model. Additionally, repeated measures correlation showed a significant correlation between the matrix and fiber viscoelastic parameters. Fourth, the results obtained from ABC demonstrated that the stochastically inferred model parameters can capture the uncertainty band of in-vitro measurements ($y_{obs}$). Fifth, the PCA results demonstrated that the posterior parameters are not separated between sites. For clinical implementation, this observation may suggest that treatment or rehabilitation strategies regarding the mechanical response of diabetic tendons may be generalized to target the entire tendon rather than focusing on specific sites. However, additional data are needed to investigate whether these findings hold in different patient populations.
Although ABC is conceptually simple and can be easily implemented for calibration purposes, it can be computationally expensive due to the high processing time in the simulator. However, we believe that this issue can be addressed by using machine or deep learning algorithms as a surrogate model instead of obtaining $y_{sim}$ observations from the simulator during each step of the ABC sampling update. Once the input parameter set $\theta_{N}$ and the corresponding observations $y_{sim}$ from the forward simulator are trained in the algorithm, unknown simulation observations based on the input parameters sampled in ABC can be predicted using the surrogate model.
Mechanical characterization of soft tissue is usually performed by testing n samples and then averaging the model parameters calibrated for n data sets. In some cases, the average of n measurements is taken first, and then non-linear curve fitting techniques for the constitutive model are applied to calibrate the parameters for a single data set. In both methods, the calibrated parameters cannot fully capture the uncertainties in the measurements. Moreover, while uncertainties arising from geometry are often not considered, the proposed method can easily incorporate these uncertainties.
This study shows that the parameters calibrated using the probabilistic technique (Table 3) deviate from those obtained through the frequentist approach (Table 2). In the frequentist method, the parameter of each specimen is estimated independently based solely on BO. In contrast, the ABC approach incorporates prior knowledge and updates the parameter estimates probabilistically. For example, consider the elastic modulus of fiber $E_{f}$ for the distal site; using the frequentist approach, the average $E_{f}$ value for nine specimens is 204.82 ± 27.42 MPa (mean ± standard deviation). However, when calibrated using ABC, the estimated $E_{f}$ is 216.2 ± 31.12 MPa. This difference highlights a fundamental distinction between the two approaches. However, the following limitations should be addressed in future studies. First, the number of samples is important for the quantification of uncertainty; hence, more specimens should be used to investigate the posterior parameters of the population. Second, investigations are performed on samples with approximately similar geometrical dimensions. Further studies can include geometrical uncertainty. Third, while the concept of ABC is simple and widely applicable, it is computationally expensive as it requires simulating the forward model many times. Additional studies can explore the use of surrogate models based on machine or deep learning algorithms.
In conclusion, the likelihood-free uncertainty quantification approach provides a useful framework for the stochastic calibration of constitutive material model parameters without the need to derive a likelihood function. This general framework can be applied to various combinations of loading data and constitutive models. In medicine, microscale characterization and uncertainty quantification of constituents that cannot be easily determined using experimental techniques, including constituent parameters, constituent stress, and their correlations, could help create more effective therapeutic strategies.
[…]
Supplementary Materials
Figure S1: Visualization the sample points for a typical specimen during Bayesian optimization. (The red star shows the best-found parameters)
Figure S2: The influence of each search space on objective function observed for a typical specimen in Bayesian optimization (The red star shows the best-found parameters
Figure S3: Posterior joint plots between parameters of the matrix and fiber component for the distal site
Figure S4: Posterior joint plots between parameters of the matrix and fiber component for the middle site
Figure S5: Posterior joint plots between parameters of the matrix and fiber component for the proximal site
Figure S6: Inferred matrix and fiber stresses for the middle and proximal sites